OpenStudy (anonymous):

Convert the equation to the standard form for a hyperbola by completing the square on x and y. x2 - y2 + 6x - 4y + 4 = 0

OpenStudy (anonymous):

(x^2)-(y^2) + 6x -4y +4 = 0 (x^2) + 6x -(y^2) - 4y + 13 = 9 ((x+3)^2) -(y^2) - 4y +4 = 9 -((x+3)^2) + (y^2) + 4y -4 = -9 -((x+3)^2) + (y^2) + 4y = -5 -((x+3)^2) + (y^2) + 4y +4 = -1 -((x+3)^2) + ((y+2)^2) = -1 ((x+3)^2) - ((y+2)^2) = 1 (x + 3)2 - (y + 2)2 = 1 is this right